Age of chetan in 2002 was 1/90 of his birthyear.wat is his age in 2006?
30
28
26
22

Answer by abdullahkhawer(5)

In these type of questions the birth year is not known. Therefore, we will have to check all answers.
-----------
Checking of Answers:
Let the birth year be x.
Let y = (1/90 * x) = chetan's age in 2002
Therefore, x = 90 * y = birth year
If y = 30 then x = 2700 > 2002. Therefore, this option is not possible.
If y = 28 then x = 2520 > 2002. Therefore, this option is not possible.
If y = 26 then x = 2340 > 2002. Therefore, this option is not possible.
If y = 22 then x = 1980 < 2002. Therefore, this option is possible.
Therefore, chetan is 22 years old in 2002.
-----------
Question:
what is his age in 2006?
As chetan is 22 years old in 2002, therefore his age in 2006 will be 22 + 4 = 26 because the difference b/w 2002 and 2006 is 4.
-----------
Ans:
Chetan is 26 years old in 2006
-----------
Summary:
You just have to check the possibility of answers using the above mentioned method. And then just add the value of difference in the value of possible answer.
-----------
Remarks:
I hope that helped you. :)